# Better Gaussian Mechanism using Correlated Noise

### v1

@@ 1 Introduction
The Gaussian mechanism is one of the most important tools in differential privacy. In its most basic form, the mechanism first computes the answer to $d$ real-valued queries. The output is then perturbed by adding noise sampled from a zero-mean Gaussian distribution independently to the answer of each query. The magnitude of the Gaussian noise required to satisfy differential privacy depends on the desired privacy parameters and the $\ell_{2}$ sensitivity of the queries. The mechanism is used as a key subroutine in a long list of more complicated differential privacy mechanisms such as DP-SGD [ACG+16] and the Census TopDown Algorithm [AAC+22] to name just two examples (The TopDown Algorithm uses the discrete variant of the mechanism [CKS20]).
@@
The Gaussian mechanism was introduced in some of the earliest work on differential privacy [DMNS06, DKM+06]. Since then the mechanism has been subject to a lot of study due to its importance. Notable results include a tight analysis under approximate differential privacy [BW18] as well as other definitions of differential privacy [BS16, Mir17, DRS22]. An adjacent line of work has improved the analysis of mechanisms where Gaussian noise is a key element. Recent results include an exact analysis of the Gaussian Sparse Histogram Mechanism [WKZK24] and development of numerical tools that allow for tighter privacy accounting used for DP-SGD [SMM19, KJH20, KJPH21, ZDW22]. These results give stronger privacy guarantees for various settings which leads to an improved privacy-utility trade-off. Improving the Gaussian mechanism and its variants by even small constant factors can have a significant impact on downstream applications. One notable such example is the accuracy of models trained with DP-SGD which can vary significantly based on the amount of noise (see e.g. [DBH+22, Figure 1]).
@@
As mentioned above, the standard Gaussian mechanism privately releases queries by adding noise with magnitude scaled to the $\ell_{2}$ sensitivity. The $\ell_{2}$ sensitivity is defined as the biggest distance in the $\ell_{2}$ norm between the query outputs for any pair of neighboring datasets. If we have no guarantees on the $\ell_{2}$ norm of the query outputs we can clip data points before adding noise. This is common practice for several tasks such as approximating gradients for private machine learning (e.g. [ACG+16, KMS+21]) and mean estimation (e.g. [BDKU20, HLY21, ALNP24]). The Gaussian mechanism can be applied to many settings because we only need to bound the $\ell_{2}$ sensitivity to privately release an estimate to any real-value queries. However, in some cases we can achieve better error guarantees by utilizing the structure of the queries.
@@
In this work we focus on settings where we want to estimate the sum of all data points in a dataset. We refer to the set of all possible differences between the non-private sum for neighboring dataset as the sensitivity space. We utilize a simple structure of the sensitivity space present for certain queries. Our technique reduces the magnitude of noise for such queries over the standard Gaussian mechanism that only considers the maximum $\ell_{2}$ norm of the sensitivity space. The structure we consider most naturally occurs under the add/remove neighboring relation. Before presenting our technique it is worth comparing the Gaussian mechanism under the add/remove and replacement neighboring relations for a query where the structure we utilize is not present. For many tasks, the $\ell_{2}$ sensitivity under the replacement neighboring relation is twice that of the add/remove relation. For example, if we clip the $\ell_{2}$ norm of data points in $\mathbb{R}^{d}$ to some value $C>0$ the $\ell_{2}$ sensitivity is clearly $C$ under the add/remove relation. But the $\ell_{2}$ sensitivity is $2C$ under replacement because we can replace any vector $x\in\mathbb{R}^{d}$ where $\|x\|_{2}=C$ with $-x$.
@@
Theorem 10 (simplified). Let $f(X)\in\mathbb{Z}^{d}$ be the answer to $d$ counting queries on dataset $X$. Define random variables such that $\eta\sim\mathcal{N}(0,\frac{\sqrt{d}+1}{4\mu^{2}})$, $Z\sim\mathcal{N}(0,\frac{d+\sqrt{d}}{4\mu^{2}}I_{d})$, and $W\sim\mathcal{N}(0,\frac{d}{\mu^{2}}I_{d})$ for any $\mu>0$. Then the mechanism that outputs $f(X)+\eta{\bf 1}^{d}+Z$ has the same differential privacy guarantees under the add/remove neighboring relation as the mechanism that outputs $f(X)+W$.
@@
Many data analysis tasks depend on the size of the input dataset $n$. However, since the size of neighboring datasets differ by $1$ under the add/remove neighboring relation we cannot release the exact size under differential privacy. This is often disregarded as a technical detail because we can estimate the size of the dataset using additional privacy budget. Some private algorithms explicitly assume that we have access to either an estimate or a bound of $n$ (see e.g. [DL09, NTZ16]). Other mechanisms implicitly rely on an estimate of the dataset size such as some versions of DP-SGD ([ACG+16]) since gradients are usually normalized based on the expected batch size which is a function of $n$. Our mechanism also relies on an estimate of the dataset size but we do not assume that an estimate is known. We instead use part of our privacy budget to compute an estimate. There is a trade-off between the accuracy of this estimate and the error for each query which we discuss in Section 3.1. If we have access to an estimate of the dataset size we can reduce the magnitude of noise added to each query from $(\sqrt{d}+1)/2$ to $\sqrt{d}/2$.
Our mechanism is simple and it is likely to be easily implementable in libraries that already support the standard Gaussian mechanism. We argue that simplicity is a desirable feature of practical differential private mechanisms. In many cases (e.g. the 2020 US Census [AAC+22]) the privately released data will be used by data analysts who might not be familiar with the theory behind differential privacy. Simple mechanisms make it easier to communicate how the data was protected and take the effect of introduced noise into account in the analysis. Furthermore, it is well known that implementations of even basic concepts and mechanisms from differential privacy requires care to avoid privacy issues (See e.g. [Mir12, CSVW22, HDH+22]). We discuss some considerations for implementation at the end of Section 3. We believe that our technique has the potential to improve mechanisms for other settings apart from counting queries. In Section 4 we adapt our core idea to a setting where we add noise with a smaller magnitude than the standard Gaussian mechanism under both the add/remove and replacement neighboring relations. We conclude our paper by discussing related work in Section 5 and potential future directions in Section 6.
@@ 2 Preliminaries
Differential Privacy. Differential privacy [DMNS06] is a statistical property of a randomized mechanism. Intuitively, a randomized mechanism satisfies differential privacy if the output distribution of the mechanism is not affected significantly by any individual. The notion of a pair of neighboring datasets, denoted as $X\sim X^{\prime}$, is used to model the impact of any one individual’s data. The two most commonly used definitions are the add/remove and replacement variants where one neighboring dataset is obtained from the other by adding, removing, or replacing any data point, respectively. We consider both definitions in this paper but focus mainly on the add/remove neighboring relation.
We present our mechanism in terms of Gaussian differential privacy ($\mu$-GDP) [DRS22]. Informally, an algorithm satisfies $\mu$-GDP if the output distributions for any neighboring datasets are at least as similar as two Gaussian distributions. This is more precisely defined in Definition 1.
Definition 1 (Gaussian differential privacy [DRS22, Definition 4]).
@@
Here $T(P,Q):[0,1]\rightarrow[0,1]$ denotes a trade-off function for two distributions $P$ and $Q$ defined on the same space. The trade-off function gives a lower bound on the type 2 error achievable for any level of type 1 error. For more technical details on trade-off functions and $f$-differential privacy we refer to Definitions 2 and 3 of [DRS22].
Note that our results do not rely on Gaussian differential privacy and can be easily translated to other popular definitions of differential privacy by replacing $\mu$ with an appropriate value.
Remark.
Let $f\colon\mathcal{U}^{\mathbb{N}}\rightarrow\mathbb{R}$ be a function with sensitivity $1$. Let $\sigma>0$ be large enough such that $f(X)+\mathcal{N}(0,\sigma^{2})$ satisfies either $(\varepsilon,\delta)$-differential privacy [BW18], $(\alpha,\varepsilon)$-RDP [Mir17], or $\rho$-zCDP [BS16]. Then all mechanisms presented in this paper satisfy the desired differential privacy guarantee if $1/\mu^{2}$ is replaced by $\sigma^{2}$.
The sensitivity of a function is a restriction on the difference in output for any pair of neighboring datasets. The $\ell_{2}$ sensitivity of a function as defined below is a bound on the distance between the output for any pair of neighboring datasets. The sensitivity space is the set of all possible differences between outputs for neighboring datasets.
Definition 2 ($\ell_{2}$ sensitivity).
@@
Definition 3 (Sensitivity space).
@@
$\{f(X)-f(X^{\prime})\in\mathbb{R}^{d}|~X\sim X^{\prime}\}.$
The magnitude of noise added by the Gaussian mechanism scales linearly in the $\ell_{2}$ sensitivity and a value based on the desired privacy guarantees. Lemma 4 gives the privacy guarantees of the Gaussian mechanism. We refer to this variant as the standard Gaussian mechanism.
Lemma 4 (The Gaussian mechanism [DRS22]).
@@
Lemma 5 (Post-processing [DRS22, Proposition 4]).
@@ 3 Algorithm
In this section we present our main contribution. The pseudo-code for our mechanism is in Algorithm 1. Throughout the section we present multiple representations of our mechanism. Each of these representations has its benefits. The error guarantees are trivial to analyze in one representation while the privacy guarantees are easier to analyze in another.
@@
The motivating problem of our work is the differentially private release of $d$ counting queries. However, we present our mechanism in the more general setting of answering $d$ queries with a real value from the interval $[0,1]$. That is, a dataset $X$ consists of $n$ data points such that $x_{i}\in[0,1]^{d}$. A pair of datasets, $X\sim X^{\prime}$, are neighboring if we can obtain $X^{\prime}$ from $X$ by either adding or removing one data point. As such, we either have that (1) $|X^{\prime}|=n-1$ and $f(X)-f(X^{\prime})\in[0,1]^{d}$ or (2) $|X^{\prime}|=n+1$ and $f(X)-f(X^{\prime})\in[-1,0]^{d}$.
The core idea of our mechanism is to take advantage of the structure of this sensitivity space. The $\ell_{2}$ sensitivity of $f$ is $\sqrt{d}$. However, notice that the vector $f(X)-f(X^{\prime})$ always has $\ell_{2}$ distance at most $\sqrt{d}/2$ to either ${\bf\frac{1}{2}}^{d}$ or $-{\bf\frac{1}{2}}^{d}$ (recall that ${\bf\frac{1}{2}}^{d}$ denotes the $d$ dimensional vector where all coordinates are $1/2$). Our mechanism spends part of the privacy budget on adding noise along ${\bf 1}^{d}$. The idea is to hide the difference between $f(X)-{\bf\frac{1}{2}}^{d}$, $f(X)$, and $f(X)+{\bf\frac{1}{2}}^{d}$. This allows us to add noise with smaller magnitude to each query. A visualization of the geometric intuition behind this idea for $d=2$ is shown in Figure 1.
Figure 1: 2D geometric intuition of our technique. Each data point is in $[0,1]\times[0,1]$. The red cross marks $f(X)$ in all figures. The two left figures depict the shape of noise for the standard Gaussian mechanism. Informally, the mechanism is private if $f(X^{\prime})$ always is inside the blue circle with radius $\sqrt{2}$. The value of $f(X^{\prime})$ under the replacement neighboring relation can be anywhere in the grey box in the first figure. That is, the box shows the result of adding any point from the sensitivity space to $f(X)$. The boxes in the second figure similarly depict the sensitivity space under the add/remove neighboring relation. In general, there are $2^{d}$ possible values for $f(X^{\prime})$ that touch the hypersphere under replacement but only $2$ such values under add/remove for any $d$. In the third figure we focus on the case where we add a data point to $X$ to construct $X^{\prime}$. Notice that the orange circle with radius 2/$\sqrt{2}/2$ centered at $f(X)+(0.5,0.5)$ contains the box. We spend part of the privacy budget to add noise along the yellow diagonal. The resulting noise is elliptical and contains the sensitivity space under add/remove as seen in the last figure.
Lemma 6.
@@
It is easy to see on Line 1 of Algorithm 1 that the error for the estimate of the $i$’th query depends only on the value of $\eta$ and $z_{i}$ for any $i\in[d]$. Since $\eta$ and $z_{i}$ are independent samples from zero-mean Gaussians their sum is distributed as a Gaussian with variance $\frac{d+\sqrt{d}}{4\mu^{2}}+\frac{\sqrt{d}+1}{4\mu^{2}}=\frac{d+2\sqrt{d}+1}{4\mu^{2}}$. Furthermore, the standard deviation is $\left(({d+2\sqrt{d}+1})/({4\mu^{2}})\right)^{1/2}=(\sqrt{d}+1)/(2\mu)$. ∎
It remains for us to prove the privacy guarantees of Algorithm 1. We first introduce another mechanism. We use a simple injective function to map each data point of $X$ to a point in $\mathbb{R}^{d+1}$. We then release an estimate of the sum of all data points in this new representation under $\mu$-GDP using the standard Gaussian mechanism. Finally, we prove that the output distribution of Algorithm 1 is equivalent to post-processing the output of this new mechanism for a fixed parameter.
@@
Fix any pair of neighboring datasets $X\sim X^{\prime}$ and let $x_{i}$ denote the data point that is either added to or removed from $X$ to obtain $X^{\prime}$. It is easy to see that $\|g(X)-g(X^{\prime})\|_{2}=\|\hat{x}_{i}\|_{2}$. By the definition of $\hat{x}_{i}$ we have that $\hat{x}_{i,j}=2\cdot x_{i,j}-1$ is in the interval $[-1,1]$ and $\hat{x}_{i,d+1}=C$. As such, it clearly holds that $\|g(X)-g(X^{\prime})\|_{2}=\sqrt{\sum_{j=1}^{d+1}\hat{x}_{i,j}^{2}}\leq\sqrt{\sum_{j=1}^{d}1^{2}+C^{2}}=\sqrt{d+C^{2}}$. ∎
Lemma 8.
@@
It follows directly from Lemmas 4 and 7. ∎
Next, we show that Algorithm 1 and the mechanism from Lemma 8 are equivalent under post-processing when setting $C=d^{1/4}$. Since post-processing does not affect the privacy guarantees, it follows directly from Lemma 8 and 9 that Algorithm 1 satisfies $\mu$-GDP.
Lemma 9.
Let $\mathcal{M}$ denote Algorithm 1 and let $\mathcal{M}^{\prime}$ denote the mechanism described in Lemma 8 with parameter $C=d^{1/4}$. Then there exists a bijection $h\colon\mathbb{R}^{d}\rightarrow\mathbb{R}^{d+1}$ such that $h(\mathcal{M}(X))$ is distributed as $\mathcal{M}^{\prime}(X)$ and $h^{-1}(\mathcal{M}^{\prime}(X))$ is distributed as $\mathcal{M}(X)$.

### v3

@@ 1 Introduction
The Gaussian mechanism is one of the most important tools in differential privacy. In its most basic form, the mechanism first computes the answer to $d$ real-valued queries. The output is then perturbed by adding noise sampled from a zero-mean Gaussian distribution independently to the answer of each query. The magnitude of the Gaussian noise required to satisfy differential privacy depends on the desired privacy parameters and the $\ell_{2}$ sensitivity of the queries. The mechanism is used as a key subroutine in a long list of more complicated differential privacy mechanisms such as DP-SGD [ACG+16] and the Census TopDown Algorithm [AAC+22] to name just two examples (The TopDown Algorithm uses the discrete variant of the Gaussian mechanism [CKS20]).
@@
The Gaussian mechanism was introduced in some of the earliest work on differential privacy [DMNS06, DKM+06]. Since then, the mechanism has been subject to a lot of study due to its importance. Notable results include a tight analysis under approximate differential privacy [BW18] as well as other definitions of differential privacy [BS16, Mir17, DRS22]. An adjacent line of work has improved the analysis of mechanisms where additive Gaussian noise is a key element. Recent results include an exact analysis of the Gaussian Sparse Histogram Mechanism [WKZK24] and development of numerical tools that allow for tighter privacy accounting used for DP-SGD (e.g. [SMM19, KJH20, KJPH21, ZDW22]). These results give stronger privacy guarantees of mechanisms for various settings, which imply an improved privacy-utility trade-off. Improving the Gaussian mechanism and its variants by even small constant factors can have a significant impact on downstream applications. One notable such example is the accuracy of models trained with DP-SGD which can vary significantly based on the amount of noise (see e.g. [DBH+22, Figure 1]).
@@
As mentioned above, the standard Gaussian mechanism privately releases queries by adding noise with magnitude scaled to the $\ell_{2}$ sensitivity. The $\ell_{2}$ sensitivity is defined as the biggest distance in the $\ell_{2}$ norm between the query outputs for any pair of neighboring datasets. If we have no guarantees on the $\ell_{2}$ norm of the query outputs we can clip data points before adding noise. This is common practice for several tasks such as approximating gradients for private machine learning (e.g. [ACG+16, KMS+21]) and mean estimation (e.g. [BDKU20, HLY21, ALNP24]). The Gaussian mechanism can be applied to many settings because we only need to bound the $\ell_{2}$ sensitivity to privately release an estimate to any real-value queries. However, in some cases we can achieve better accuracy by utilizing the structure of the queries.
@@
In this work we focus on settings where we want to estimate the sum of all data points in a dataset. We refer to the set of all possible differences between the non-private sums for neighboring dataset as the sensitivity space. We utilize a simple structure of the sensitivity space present for certain queries. Our technique reduces the magnitude of noise for such queries over the standard Gaussian mechanism that only considers the maximum $\ell_{2}$ norm of the sensitivity space. The structure we consider most naturally occurs under the add/remove neighboring relation.
Before presenting our technique it is worth considering a query where the structure we utilize is not present. We compare the sensitivity under the add/remove and replacement neighboring relations. For many tasks, the $\ell_{2}$ sensitivity under the replacement neighboring relation is twice that of the add/remove relation. For example, if we clip the $\ell_{2}$ norm of data points in $\mathbb{R}^{d}$ to some value $C>0$ the $\ell_{2}$ sensitivity is clearly $C$ under the add/remove relation. But the $\ell_{2}$ sensitivity is $2C$ under replacement because we can replace any vector $x\in\mathbb{R}^{d}$ where $\|x\|_{2}=C$ with $-x$. The magnitude of noise added by the standard Gaussian mechanism therefore differs by a factor of 2 between the two neighboring relations.
@@
Theorem 3.1 (simplified). Let $f(X)\in\mathbb{N}^{d}$ be the answer to $d$ counting queries on dataset $X$. Define random variables such that $\eta\sim\mathcal{N}(0,\frac{\sqrt{d}+1}{4\mu^{2}})$, $Z\sim\mathcal{N}(0,\frac{d+\sqrt{d}}{4\mu^{2}}I_{d})$, and $W\sim\mathcal{N}(0,\frac{d}{\mu^{2}}I_{d})$ for any $\mu>0$. Then the mechanism that outputs $f(X)+\eta{\bf 1}^{d}+Z$ has the same differential privacy guarantees under the add/remove neighboring relation as the mechanism that outputs $f(X)+W$.
@@
Many data analysis tasks depend on the size of the input dataset $n$. However, since the size of neighboring datasets differ by $1$ under the add/remove neighboring relation we cannot release the exact size under differential privacy. This is often disregarded as a technical detail because we can estimate the size of the dataset using additional privacy budget. Some private algorithms explicitly assume that we have access to either an estimate or a bound of $n$ (see e.g. [DL09, NTZ16]). Other mechanisms implicitly rely on an estimate of the dataset size such as some versions of DP-SGD [ACG+16] since gradients are usually normalized based on the expected batch size which is a function of $n$. Our mechanism also relies on an estimate of the dataset size but we do not assume that an estimate is known. We instead use part of our privacy budget to compute an estimate. There is a trade-off between the accuracy of this estimate and the error for each query which we discuss in Section 3.1. If we have access to an estimate of the dataset size we can reduce the magnitude of noise added to each query from $(\sqrt{d}+1)/2$ to $\sqrt{d}/2$.
Our mechanism is simple and it is likely to be easily implementable in libraries that already support the standard Gaussian mechanism. We argue that simplicity is a desirable feature of practical differentially private mechanisms. In many cases (e.g. the 2020 US Census [AAC+22]) the privately released data will be used by data analysts who might not be familiar with the theory behind differential privacy. It is easier to communicate how the data was protected when the differentially private mechanism is simple. It is also easier to take the effect of introduced noise into account for any downstream analysis. Furthermore, it is well known that implementations of even basic concepts and mechanisms from differential privacy requires care to avoid privacy issues (See e.g. [Mir12, CSVW22, HDH+22]). We discuss some considerations for implementation at the end of Section 3. We believe that our technique has the potential to improve mechanisms for other settings apart from counting queries. In Section 4 we adapt our core idea to a setting where we add noise with a smaller magnitude than the standard Gaussian mechanism under both the add/remove and replacement neighboring relations. We conclude our paper by discussing related work in Section 5 and potential future directions in Section 6.
@@ 2 Preliminaries
Differential Privacy. Differential privacy [DMNS06] is a statistical property of a randomized mechanism. Intuitively, a randomized mechanism satisfies differential privacy if the output distribution of the mechanism is not affected significantly by any individual. The notion of neighboring datasets is used to model the impact of any one individual’s data. The two most commonly used definitions are the add/remove and replacement variants defined below. We consider both definitions in this paper but focus mainly on the add/remove neighboring relation.
Definition 2.1 (Neighboring datasets (denoted $X\sim X^{\prime}$))
Let $X=(x_{1},x_{2},\dots,x_{n})$ be a dataset of size $n$. If $X$ and $X^{\prime}$ are neighboring datasets under the add/remove neighboring relation then there exists an $i$ such that either $X^{\prime}=(x_{1},\dots,x_{i-1},x_{i+1},\dots,x_{n})$ or $X=(x^{\prime}_{1},\dots,x^{\prime}_{i-1},x^{\prime}_{i+1},\dots,x_{n+1})$. If $X$ and $X^{\prime}$ are neighboring datasets under the replacement neighboring relation we have $|X|=|X^{\prime}|$ and there exists an $i$ such that $x_{j}=x^{\prime}_{j}$ for all $j\neq i$.
We present our mechanism in terms of Gaussian differential privacy ($\mu$-GDP) [DRS22]. Informally, an algorithm satisfies $\mu$-GDP if the output distributions for any neighboring datasets are at least as similar as two Gaussian distributions. This is more precisely defined in Definition 2.2. Note that none of our results relies on Gaussian differential privacy. In the remark below we discuss how to easily translate our results to some other popular definitions of differential privacy by replacing $\mu$ with an appropriate value.
Definition 2.2 (Gaussian differential privacy [DRS22, Definition 4])
@@
Here $T(P,Q):[0,1]\rightarrow[0,1]$ denotes a trade-off function as defined below, and the inequality in Definition 2.2 must hold for all inputs to the functions (any $0\leq\alpha\leq 1$). The trade-off function gives a lower bound on the type II error achievable for any level of type I error. We note that understanding the technical details of the trade-off function and $\mu$-GDP is not required to follow the proofs in this paper. For more background on trade-off functions and $f$-differential privacy, we refer interested readers to Definitions 2 and 3 of [DRS22].
Definition 2.3 (Trade-off function)
Let $P$ and $Q$ be any two probability distributions defined on the same space. The trade-off function $T(P,Q):[0,1]\rightarrow[0,1]$ is defined as
$T(P,Q)(\alpha)=\inf\{\beta_{\phi}:\alpha_{\phi}\leq\alpha\}\,,$
where the infimum is taken over all (measurable) rejection rules. For a rejection rule $\phi$, the values of $\alpha_{\phi}$ and $\beta_{\phi}$ denotes the type I and type II error rates, respectively.
Remark
Let $f\colon\mathcal{U}^{\mathbb{N}}\rightarrow\mathbb{R}$ be a function with sensitivity $1$. Let $\sigma>0$ be large enough such that $f(X)+\mathcal{N}(0,\sigma^{2})$ satisfies either $(\varepsilon,\delta)$-differential privacy [BW18, Theorem 8], $(\alpha,\varepsilon)$-RDP [Mir17, Corollary 3], or $\rho$-zCDP [BS16, Proposition 1.6]. Then all mechanisms presented in this paper satisfy the desired differential privacy guarantee if $1/\mu^{2}$ is replaced by $\sigma^{2}$.
The $\ell_{2}$ sensitivity of a function as defined below is a bound on the distance between the output for any pair of neighboring datasets. More generally, the sensitivity space is the set of all possible differences between outputs for neighboring datasets. The $\ell_{2}$ sensitivity equals the maximum $\ell_{2}$ norm for any element in the sensitivity space.
Definition 2.4 ($\ell_{2}$ sensitivity)
@@
Definition 2.5 (Sensitivity space)
@@
$\{f(X)-f(X^{\prime})\in\mathbb{R}^{d}|X\sim X^{\prime}\}\,.$
The magnitude of noise added by the Gaussian mechanism scales linearly in the $\ell_{2}$ sensitivity and a value based on the desired privacy guarantees. Lemma 2.1 gives the privacy guarantees of the Gaussian mechanism. We refer to this variant as the standard Gaussian mechanism throughout the paper.
Lemma 2.1 (The Gaussian mechanism [DRS22])
@@
Lemma 2.2 (Post-processing [DRS22, Proposition 4])
@@ 3 Algorithm
In this section we present our main contribution. The pseudo-code for our mechanism is in Algorithm 1. Throughout the section we present multiple equivalent representations of our mechanism. Each of these representations has its benefits. The error guarantees are trivial to analyze in first representation while the privacy guarantees are easier to analyze in the second. We present a third representation of the mechanism in Corollary 3.1, and finally in Section 3.1 we use yet another representation when discussing the trade-off between two sources of error. It is sometimes subjective which representation is preferred for a specific task. We hope that by presenting our work in multiple ways, the core idea behind our technique will become more intuitive to readers with various backgrounds.
@@
The motivating problem of our work is to release $d$ counting queries under differential privacy. However, we present our mechanism in the more general setting of answering $d$ queries with a real value from the interval $[0,1]$. That is, a dataset $X$ consists of $n$ data points such that $x_{i}\in[0,1]^{d}$. A pair of datasets, $X\sim X^{\prime}$, are neighboring if we can obtain $X^{\prime}$ from $X$ by either adding or removing one data point. As such, we either have that (1) $|X^{\prime}|=n-1$ and $f(X)-f(X^{\prime})\in[0,1]^{d}$ or (2) $|X^{\prime}|=n+1$ and $f(X)-f(X^{\prime})\in[-1,0]^{d}$.
The core idea of our mechanism is to take advantage of the structure of this sensitivity space. The $\ell_{2}$ sensitivity of $f$ is $\sqrt{d}$. However, notice that the vector $f(X)-f(X^{\prime})$ always has $\ell_{2}$ distance at most $\sqrt{d}/2$ to either ${\bf\frac{1}{2}}^{d}$ or $-{\bf\frac{1}{2}}^{d}$ (recall that ${\bf\frac{1}{2}}^{d}$ denotes the $d$ dimensional vector where all coordinates are $1/2$). Our mechanism spends part of the privacy budget on adding noise along ${\bf 1}^{d}$. The idea is to hide the difference between $f(X)-{\bf\frac{1}{2}}^{d}$, $f(X)$, and $f(X)+{\bf\frac{1}{2}}^{d}$. This allows us to add noise with smaller magnitude to each query. A visualization of the geometric intuition behind this idea for $d=2$ is shown in Figure 1. Next, we discuss the properties of our mechanism.
Figure 1: 2D geometric intuition of our technique. Each data point is in $[0,1]\times[0,1]$. The red cross marks $f(X)$ in all figures. The two left figures depict the shape of noise for the standard Gaussian mechanism. Informally, the mechanism is private if $f(X^{\prime})$ always is inside the blue circle with radius $\sqrt{2}$. The value of $f(X^{\prime})$ under the replacement neighboring relation can be anywhere in the grey box in the first figure. That is, the box shows the result of adding any point from the sensitivity space to $f(X)$. The boxes in the second figure similarly depict the sensitivity space under the add/remove neighboring relation. In general, there are $2^{d}$ possible values for $f(X^{\prime})$ that touch the hypersphere under replacement, but only $2$ such values under add/remove for any $d$. In the third figure we focus on the case where we add a data point to $X$ to construct $X^{\prime}$. Notice that the orange circle with radius 2/$\sqrt{2}/2$ centered at $f(X)+(0.5,0.5)$ contains the box. We spend part of the privacy budget to add noise along the yellow diagonal. The resulting noise is elliptical and contains the sensitivity space under add/remove as seen in the last figure.
Lemma 3.1
@@
    It is easy to see on Line 1 of Algorithm 1 that the error for the estimate of the $i$’th query depends only on the value of $\eta$ and $z_{i}$ for any $i\in[d]$. Since $\eta$ and $z_{i}$ are independent samples from zero-mean Gaussians their sum is distributed as a zero-mean Gaussian with variance $\frac{d+\sqrt{d}}{4\mu^{2}}+\frac{\sqrt{d}+1}{4\mu^{2}}=\frac{d+2\sqrt{d}+1}{4\mu^{2}}$. Furthermore, the standard deviation is $\left(({d+2\sqrt{d}+1})/({4\mu^{2}})\right)^{1/2}=(\sqrt{d}+1)/(2\mu)$.
It remains for us to prove the privacy guarantees of Algorithm 1. To do so, we first introduce another mechanism. We use a simple injective function to map each data point of $X$ to a point in $\mathbb{R}^{d+1}$. We then release an estimate of the sum of all data points in this new representation under $\mu$-GDP using the standard Gaussian mechanism. Finally, we prove that Algorithm 1 is equivalent to post-processing the output of this new mechanism for a fixed parameter.
@@
    Fix any pair of neighboring datasets $X\sim X^{\prime}$ and let $x_{i}$ denote the data point that is either added to or removed from $X$ to obtain $X^{\prime}$. It is easy to see that either $g(X)-g(X^{\prime})=\hat{x}_{i}$ or $g(X^{\prime})-g(X)=\hat{x}_{i}$. By the definition of $\hat{x}_{i}$ we have that $\hat{x}_{i,j}$ is in the interval $[-1,1]$ for any $j\in[d]$. The $\ell_{2}$ sensitivity of $g$ is thus bounded by $\|g(X)-g(X^{\prime})\|_{2}=\|\hat{x}_{i}\|_{2}=\sqrt{\sum_{j=1}^{d+1}\hat{x}_{i,j}^{2}}=\sqrt{\sum_{j=1}^{d}(2\cdot x_{i,j}-1)^{2}+C^{2}}\leq\sqrt{\sum_{j=1}^{d}1^{2}+C^{2}}=\sqrt{d+C^{2}}$.
Lemma 3.3
@@
    It follows directly from Lemmas 2.1 and 3.2.
Next, we show that Algorithm 1 and the mechanism from Lemma 3.3 are equivalent under post-processing when setting $C=d^{1/4}$. We set $C=d^{1/4}$ because that choice of $C$ minimizes the error for each query. We discuss the effect of using other values for the parameter in more detail in Section 3.1.
We recover a private estimate of $f(X)$ from the private estimate of $g(X)$ by inverting the transformation we performed when constructing $\hat{X}$ from $X$. Notice that $f(X)_{i}=(g(X)_{i}-g(X)_{d+1}/C)/2$ which allows us to estimate $f(X)_{i}$ using the private estimates of $g(X)_{i}$ and $g(X)_{d+1}$. Since post-processing does not affect the privacy guarantees (Lemma 2.2), it follows directly from Lemma 3.3 and 3.4 that Algorithm 1 satisfies $\mu$-GDP.
Lemma 3.4
Let $\mathcal{M}$ denote Algorithm 1 and let $\mathcal{M}^{\prime}$ denote the mechanism described in Lemma 3.3 with parameter $C=d^{1/4}$. Then there exists a bijection $h\colon\mathbb{R}^{d}\rightarrow\mathbb{R}^{d+1}$ such that $h(\mathcal{M}(X))$ is distributed as $\mathcal{M}^{\prime}(X)$ and $h^{-1}(\mathcal{M}^{\prime}(X))$ is distributed as $\mathcal{M}(X)$.